OpenStudy (anonymous):

Divide and simplify \[\frac{ 1 }{ 4 }div \frac{ 1 }{ 36 }\]

OpenStudy (anonymous):

\[\frac{ 1 }{ 8}\] simplify it down to \[\frac{ 1 }{ 4 }\]and than simplify that to \[\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

your welcome

OpenStudy (texaschic101):

when you are dividing fractions, you " flip " what you are dividing by, then multiply. But make sure you are flipping the right number. And if you can reduce the fractions beforehand, then do so.

OpenStudy (texaschic101):

example : 1/2 divided by 3/4 = 1/2 * 4/3 = 4/6 which reduces to 2/3. do not do it this way... 1/2 divided by 3/4 = 2/1 * 3/4 = 6/4 (this is wrong...flip what you are dividing by)

OpenStudy (anonymous):

thanks and i kno the reciprocal

OpenStudy (texaschic101):

1/4 divided by 1/36 = 1/4 * 36/1 = 36/4 which reduces to 9

OpenStudy (anonymous):

so 1/2 is incorrect ?

OpenStudy (texaschic101):

yes

OpenStudy (anonymous):

ah okay . thanks

OpenStudy (texaschic101):

sure thing :)